Which of the following pairs are NOT isoelectronic?

N3- and P3-

S2- and Cl-

F- and Mg2+

Na+ and O2-
Answer #1

Isoelectronic means species which have same number of electrons

N3- 7+3 = 10 electrons and P3- 15+3 = 18 electron it is not isoelectronic

S2- 16+2 = 18 electron and Cl- 17+1 = 18 electron it is isoelectronic

F- 9+1 = 10 electron and Mg2+ 12-2 = 10 electron it is isoelectronic

Na+ 11-1 = 10 electron and O2- 8+2 = 10 electron it is isoelectronic

So the correct answer is option (A) N3- and P3-

Please give me thumbs up